# Is Human-Like Text Liked by Humans? Multilingual Human Detection and Preference Against AI

### v1

@@ D.1 Arabic
These adjustments addressed critical gaps appeared in prior outputs. The newly-generated tweets touched on relatable human topics and expressed genuine emotions tied to daily experiences. We used GPT-4o for generation as it yielded promising results, while Qwen is underperformed. As mentioned in Appendix C.1, Qwen texts often include non-Arabic words within the tweets. To evaluate the effectiveness of the improved prompts, 200 sampled tweets were assessed by annotators, determining whether the outputs addressed the identified gaps. The evaluation revealed that 53% of the tweets fully met the criteria, 29% partially did, and 18% did not. However, some limitations were observed: GPT-4o frequently added irrelevant hashtags at the end of tweets, making the outputs identifiable as machine-generated, and the tweets often carried an overly optimistic tone. Even when negative experiences were mentioned, the sentiment leaned toward positivity and new beginnings. Overall, while these improved prompts succeeded in filling some previous gaps, new issues emerged, which could be addressed with more precise instructions, such as explicitly avoiding hashtags or adopting a more somber tone when required.
@@
Building on the identified discrepancies between human-written and machine-generated summaries discussed in Appendix C.1, the generation prompt was refined to address these gaps, resulting in the following prompts:
@@
Before prompt engineering, the accuracy of annotating which of two texts is machine or human written was 100%. After prompt engineering, the accuracy dropped to 66% indicating that prompt engineering works at least partially to bridge the gap between the writing styles of humans and machines. We follow the same model choices and data subset outlined in B.1 but we enhance the prompt as shown below:
@@ D.2 Chinese
D.3 English Peer Meta Review
@@ D.4 Hindi News
D.5 Italian DICE News
@@
D.6 Japanese News
Based on the human detection evaluation, we found several distinctive stylistic characteristics in LLM-generated Japanese news with the original prompt, as detailed in C.6. To prevent LLMs from generating such easily detectable clues, we provide the following improved prompt:
@@ D.9 Vietnamese
D.10 Subjectivity in Fill-the-gap Survey
@@
D.11 Automatic Detection Data and Results
Table 9 presents the distribution of data used in automatic MGT detection, and the results are shown in Figure 6. 19 among 26 automatic detection approaches demonstrate lower detection accuracy on the improved generations produced by the adjusted prompts. Refer to caption Figure 6: Detection accuracy differences of 26 automatic machine-generated text detection approaches on original vs. improved generations. Source / Domain Language # Improved # Original LLM Generator List QA Chinese 3422 9842 GPT-4o (3421), GPT-4o-mini (6845), GPT-4o-2024-05-13 (2998) Essay Chinese 849 702 Claude-3-5-Sonnet (773), GLM-4-9B-Chat (778) GovReport Chinese 16776 0 Baichuan2-13B-Chat (5521), ChatGLM3-6B (5359), GPT-4o (5896) News Hindi 0 1199 GPT-4 (600), Human (599) News Japanese 0 300 GPT-4o (300) News Russian 5915 600 GPT-4o (3921), Vikhrmodels/Vikhr-Nemo-12B-Instruct-R-21-09-24 (3224) Sunmary Arabic 153 153 GPT-4o (306) Sunmary Russian 5944 585 Vikhrmodels/Vikhr-Nemo-12B-Instruct-R-21-09-24 (3279), GPT-4o (3300) Tweets Arabic 2800 214 GPT-4-Turbo (1400), GPT-4 (1545), Qwen2.5 72B (69) Total 32,487 17,017 Table 9: Statistics of data used in automatic detection: generations using original and the improved prompts.
D.12 Why do Human Preference strongly differ in Zhihu QA?

### v2

@@ E.1 Arabic
These adjustments addressed critical gaps appeared in prior outputs. The newly-generated tweets touched on relatable human topics and expressed genuine emotions tied to daily experiences. We used GPT-4o for generation as it yielded promising results, while Qwen is underperformed. As mentioned in Appendix D.1, Qwen texts often include non-Arabic words within the tweets. To evaluate the effectiveness of the improved prompts, 200 sampled tweets were assessed by annotators, determining whether the outputs addressed the identified gaps. The evaluation revealed that 53% of the tweets fully met the criteria, 29% partially did, and 18% did not. However, some limitations were observed: GPT-4o frequently added irrelevant hashtags at the end of tweets, making the outputs identifiable as machine-generated, and the tweets often carried an overly optimistic tone. Even when negative experiences were mentioned, the sentiment leaned toward positivity and new beginnings. Overall, while these improved prompts succeeded in filling some previous gaps, new issues emerged, which could be addressed with more precise instructions, such as explicitly avoiding hashtags or adopting a more somber tone when required.
@@
Building on the identified discrepancies between human-written and machine-generated summaries discussed in Appendix D.1, the generation prompt was refined to address these gaps, resulting in the following prompts:
@@
Before prompt engineering, the accuracy of annotating which of two texts is machine or human written was 100%. After prompt engineering, the accuracy dropped to 66% indicating that prompt engineering works at least partially to bridge the gap between the writing styles of humans and machines. We follow the same model choices and data subset outlined in C.1 but we enhance the prompt as shown below:
@@ E.2 Chinese
E.3 English Peer Meta Review
@@ E.4 Hindi News
E.5 Italian DICE News
@@
E.6 Japanese News
Based on the human detection evaluation, we found several distinctive stylistic characteristics in LLM-generated Japanese news with the original prompt, as detailed in D.6. To prevent LLMs from generating such easily detectable clues, we provide the following improved prompt:
@@ E.9 Vietnamese
E.10 Subjectivity in Fill-the-gap Survey
@@
Appendix F Automatic Detection Data and Results
Table 9 presents the distribution of data used in automatic MGT detection, and the results are shown in Figure 6. 19 among 26 automatic detection approaches demonstrate lower detection accuracy on the improved generations produced by the adjusted prompts.
Figure 6: Detection accuracy differences of 26 automatic machine-generated text detection approaches on original vs. improved generations.
Source / Domain Language # Improved # Original LLM Generator List
QA Chinese 3422 9842 GPT-4o (3421), GPT-4o-mini (6845), GPT-4o-2024-05-13 (2998)
Essay Chinese 849 702 Claude-3-5-Sonnet (773), GLM-4-9B-Chat (778)
GovReport Chinese 16776 0 Baichuan2-13B-Chat (5521), ChatGLM3-6B (5359), GPT-4o (5896)
News Hindi 0 1199 GPT-4 (600), Human (599)
News Japanese 0 300 GPT-4o (300)
News Russian 5915 600 GPT-4o (3921), Vikhrmodels/Vikhr-Nemo-12B-Instruct-R-21-09-24 (3224)
Sunmary Arabic 153 153 GPT-4o (306)
Sunmary Russian 5944 585 Vikhrmodels/Vikhr-Nemo-12B-Instruct-R-21-09-24 (3279), GPT-4o (3300)
Tweets Arabic 2800 214 GPT-4-Turbo (1400), GPT-4 (1545), Qwen2.5 72B (69)
Total 32,487 17,017
Table 9: Statistics of data used in automatic detection: generations using original and the improved prompts.
Appendix G Human Preference
G.1 Why do Human Preference Strongly Differ in Zhihu QA?
@@
G.2 Preference Distributions Vary Across Individuals.
The annotators for Russian and Arabic exhibited similar preference distributions, whereas large differences occurred for Chinese QA. For example, in the Chinese Zhihu QA, the second annotator selected only seven human-written texts, while the fourth one chose 284 (95%). In the QA-emo, the first and second annotators preferred human-written responses, whereas the third and the fourth favored machine-generated text. This variance shows the charm of collecting personal preferences and then optimizing models to align with individual philosophies. Our preference annotations can serve as a valuable resource for investigating the relationship between annotator characteristics (e.g., MBTI personality, gender, and age) and preferences. Also, the data can guide models to match individual preferences in multilingual contexts.